OpenStudy (anonymous):

Find the x-intercept and the y-intercept for the graph of the equation. x+3y=-9 a. –10; –2 c. –9; –4 b. –9; –3 d. –8; –3

OpenStudy (anonymous):

the first thing you would do is to get the y on one side, then the next step is to divide the -9 with the 3 to get the x intercept.

OpenStudy (texaschic101):

there is an easier way

OpenStudy (texaschic101):

x + 3y = -9 To find x intercept, sub in 0 for y and solve for x. To find y intercept, sub in 0 for x and solve for y.

OpenStudy (anonymous):

i really do not understand

OpenStudy (texaschic101):

I will do one... x intercept (sub in 0 for y) x + 3(0) = -9 x + 0 = -9 x = -9 ....so the x intercept is (-9,0) can you find the y intercept by subbing in 0 for x ??

OpenStudy (texaschic101):

Understand ?

OpenStudy (texaschic101):

the way your answer choices are worded, the x intercept is -9 to find y intercept, sub in 0 for x x + 3y = -9 0 + 3y = -9 3y = -9 -- divide both sides by 3 y = -9/3 y = -3 y intercept is -3 so your solution is (-9.-3)

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

No, i understand but thank you.

OpenStudy (texaschic101):

sure thing :)

OpenStudy (anonymous):

wait i need help on another

OpenStudy (anonymous):

3x+y=-3

OpenStudy (anonymous):

a. 0; –3 c. –1; –4 b. –1; –3 d. –2; –2

OpenStudy (texaschic101):

3x + y = 3 find the x intercept by subbing in 0 for y and solving for x. Try it

OpenStudy (texaschic101):

oops...3x + y = -3

OpenStudy (texaschic101):

you can do this

OpenStudy (texaschic101):

let me start you off... Looking for x intercept..sub in 0 for y 3x + 0 = -3 can you finish ?

OpenStudy (anonymous):

do i divide next ?

OpenStudy (texaschic101):

3x = -3 --- divide both sides by 3

OpenStudy (anonymous):

so it would be -1;-3?

OpenStudy (texaschic101):

3x/3 = -3/3 ---(the 3's on the left side cancel out) x = -1 (-3/3 = -1) x intercept is -1 3x + y = -3 y intercept, sub in 0 for x 3(0) + y = -3 y = -3 yes....solution is (-1,-3)

OpenStudy (anonymous):

thank you for all your help (:

OpenStudy (texaschic101):

even if your equation is in slope intercept form, y = mx + b, you can still find the intercepts the same way.

OpenStudy (anonymous):

what is a slope?

OpenStudy (texaschic101):

http://www.basic-mathematics.com/what-is-slope.html ok......to find the slope, you need to put the equation in y = mx + b form, and the slope is in the m position. For example, 2x + 4y = 9 -- subtract 2x from both sides 4y = -2x + 9 -- divide both sides by 4 y = -2/4x + 9/4 -- reduce anything that reduces y = -1/2x + 9/4 The number in the m position is the slope. Therefore, the slope of this equation is -1/2